A company's profit function, in dollars, is given by P(x)= -0.3x^2+56.16x-638 where x is the number of items sold. in addition, the company has a marginal revenue function of MR(x)= -0.6x+76.16. determine the total cost for the company to produce 10 items .


1
Expert's answer
2021-06-02T16:50:36-0400

since, profit=revenue-cost

This implies,

cost=revenue-profit

Finding revenue function from the given marginal revenue function,

"R(x)=\int MR(x)dx=\int (-0.6x+76.16)dx=
-0.3x^2+76.16x+c\newline
\text{When items is zero i.e., }x=0, \newline then
R(0)=0\implies c=0.\newline
R(x)=-0.3x^2+76.16x"

Therefore,

"C(x)=R(x)-P(x)=
(-0.3x^2+76.16x)+(-0.3x^2+56.16x-638)=-0.3x^2+132.32x-638\newline
C(10)=-0.3(10)^2+132.32(10)-638
=655.2\newline
\text{tuhs, total cost for the company to produce 10 items is 655.2.}"